$SPAD/src/input f04mcf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)cl all
showArrayValues true
showScalarValues true
n:=6
al:Matrix SF:=
 [[1.0 ,2.0 ,1.0 ,3.0 ,1.0 ,1.0 ,5.0 ,4.0 ,1.5 ,0.5 ,1.0 ,1.5 ,5.0,1.0 ]]
lal:=14
d:Matrix SF:=
 [[1.0 ,1.0 ,4.0 ,16.0 ,1.0 ,16.0 ]]
nrow:Matrix Integer:=
 [[1 ,2 ,2 ,1 ,5 ,3 ]]
ir:=2
b:Matrix SF:=
 [[6 ,-10 ],_
  [15 ,-21 ],_
  [11 ,-3 ],_
  [0 ,24 ],_
  [51 ,-39 ],_
  [46 ,67 ]]
nrb:=6
iselct:=1
nrx:=6
result:=f04mcf(n,al,lal,d,nrow,ir,b,nrb,iselct,nrx,-1)

———-

References^{[1] nothing}

4